OpenStudy (anonymous):

How many photons are produced in a laser pulse of 0.819 J at 531 nm?

OpenStudy (aaronq):

find the energy given by one photon of that wavelength E=hc/lambda divide the energy given 0.819 J by the energy of one photon

OpenStudy (anonymous):

\[3.74 \times 10^-37\]?

OpenStudy (aaronq):

what energy did you get per photon

OpenStudy (aaronq):

0.000000000374 = 3.74*10^-10 J /photon

OpenStudy (anonymous):

Divide that that by .819?

OpenStudy (aaronq):

the other way around 0.819 J/3.74*10^-10 J = # of photons

OpenStudy (anonymous):

Ah.\[2.19\times10^9\]

OpenStudy (aaronq):

that sounds right